OpenStudy (anonymous):

The expression 4x-3 will equal 5 when x equals which of the following? 1 2 3 4 5

OpenStudy (anonymous):

2

OpenStudy (anonymous):

4x-3=5 4x-3+3=5+3 4x=8 x=2